OpenStudy (anonymous):

Please Help Last One!! A dispenser holds of 5500 cm3 of liquid soap and is now full. The radius of the dispenser is 7.5 cm. What is the difference between the height of the soap in the full dispenser and the height when 3355 cm3 of soap remains in the dispenser? Use 3.14 to approximate pi, and round your answer to one decimal place.

OpenStudy (anonymous):

@iGreen. @sleepyhead314

OpenStudy (anonymous):

Um..does it say what shape the dispenser is..?

OpenStudy (anonymous):

a cylinder

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

Formula: \(V = \pi r^2 h\) Plug in what we know: \(5500 = (3.14)(7.5^2)(h)\) Simplify & solve for the height

OpenStudy (anonymous):

176.625

OpenStudy (anonymous):

Yes, now divide that to 5500 to find the height.

OpenStudy (anonymous):

i got 0.0321136363636364

OpenStudy (anonymous):

Not 176.625 / 5500, 5500 / 176.625

OpenStudy (anonymous):

oh i got 31

OpenStudy (anonymous):

Yes, now we do: \(3355 = (3.14)(7.5^2)(h)\) Solve for 'h'

OpenStudy (anonymous):

how do we do that

OpenStudy (anonymous):

Simplify \((3.14)(7.5^2)\) first.

OpenStudy (anonymous):

176.625

OpenStudy (anonymous):

Yes, now we have: \(3355 = 176.625h\) Divide 176.625 to both sides

OpenStudy (anonymous):

to both sides of what??

OpenStudy (anonymous):

The equation. \(\dfrac{3355}{176.625} = \dfrac{176.625h}{176.625}\) 176.625 cancels on the right side, so just divide 3355 / 176.625

OpenStudy (anonymous):

i got 18.

OpenStudy (anonymous):

Yes, so now subtract 31 - 18

OpenStudy (anonymous):

13

OpenStudy (anonymous):

Actually it should be: 31.1394197 - 18.995046

OpenStudy (anonymous):

Those are exact..and you should've rounded to 19 not 18 :P

OpenStudy (anonymous):

i got 12.1443737

OpenStudy (anonymous):

so the answer is 12

OpenStudy (anonymous):

12.1 It says to round to 1 decimal point..

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

Thnx so much